Joe's starting salary as a mechanical engineer is around $70,000. Joe is planning to place a total of 12​% of his salary each year in the mutual fund. Joe expects a 3​% salary increase each year for the next 25 years of employment. If the mutual fund will average 7​% annual return over the course of his​ career, what can Joe expect at​ retirement?

The amount Joe can expect at retirement is $__________ thousand. ​(Round to the nearest whole​ number.)
Answer #1

Salary = 70000

Payment at EOY 1 = 70000*12%/100 = 8400

P = 8400*[1-(1+G)^N(1+I)^-N/I-G]

= 8400[1-(1+0.03)^25(1+0.07)^-25/0.07-0.03]

= 8400*[1-(2.09378)(0.18425)/0.04]

= 8400*[1-0.38578/0.04]

= 8400*15.3555

= 128986.2

The future worth = 128986.2*(1+0.07)^25

= 128986.2*5.42743

= 700064